Bramble Corp. purchased equipment in January of 2008 for $412000. The equipment was being depreciated on the straight-line method over an estimated useful life of 20 years, with no salvage value. At the beginning of 2018, when the equipment had been in use for 10 years, the company paid $55000 to overhaul the equipment. As a result of this improvement, the company estimated that the useful life of the equipment would be extended an additional 5 years. What should be the depreciation expense recorded for this equipment in 2018? $17400 $20600 O $13500 O $10300
Answer #1
Carrying value = 412000*10yr/20yr = 206000
Remaining life = 15 i.e. 10+5
Depreciable cost = 206000+55000 = 261000
Depreciation expenses = Depreciable cost/ remaining life
Depreciation exp = 261000/15 = 17400
Option $ 17400 is CORRECT